OpenStudy (anonymous):

Can someone help me please ?! What are the solutions of the equation Z^2-6z-27=0

OpenStudy (anonymous):

1. 3,9 2. 3,-9 3.-3,9 4.-3,-9

OpenStudy (irishboy123):

the solution to \[a z^2 + bz + c = 0\] is\[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] plug and play.

OpenStudy (zehanz):

Or, try to factor \(z^2-6z-27\), that is, write it as a product like this: \((z-a)\cdot(z-b)=0\). If you can do this, you can split the equation into two easier ones: \(z-a=0\) and \(z-b=0\), because if one factor of a product is zero, than the whole product is. In this case you can see the solutions are z=a or z=b. The only question is: how to find a and b? Well, they are not that hard to find: The sum of a and b must be -6 (the coefficient of z), so \(a+b=-6\). The product of a and b must be -27 (the constant), so \(a \cdot b=-27\). If you think a while about it, you will be able to find a and b. My guess is, you already know about this method...

OpenStudy (zehanz):

So the equation can be rewritten as \((z-9)\cdot (z+3)=0\), see: \(-9 + 3 = -6\) and \(-9 \cdot 3=-27\). Can you now see what the right answer is?

OpenStudy (anonymous):

Thanks guys

OpenStudy (zehanz):

YW!